OpenStudy (anonymous):

Evaluate -9 - (-4) 5 -5 13 -13

OpenStudy (jiteshmeghwal9):

-*-=+

OpenStudy (anonymous):

on my test i chose the answer - 13 but it is saying it is wrong

OpenStudy (anonymous):

- ( -) = +

OpenStudy (anonymous):

-(-4)=?

OpenStudy (jiteshmeghwal9):

-(-4)=?

OpenStudy (anonymous):

can someone please explain

OpenStudy (anonymous):

- 9 -(-4) = -9 + 4= -5

OpenStudy (anonymous):

two negatives cancels out when you multiply or divide them

OpenStudy (anonymous):

- ( -) = +

OpenStudy (anonymous):

cancel*

OpenStudy (kaederfds):

-9+4 = -5

OpenStudy (anonymous):

I don't have nothing to do with this..

OpenStudy (lgbasallote):

why is everyone repeating the same things?

OpenStudy (anonymous):

but this is not multiplication or division isint is subtraction and addition

OpenStudy (lgbasallote):

i agree @mumina1999

OpenStudy (anonymous):

the parenthesis is MULTIPLICATION

OpenStudy (anonymous):

i dont think so the parentheses are just there to show that 4 is a negative number

OpenStudy (lgbasallote):

the parentheses are there to make - -4 prettier

OpenStudy (anonymous):

yup what he/ she said ^^^^^

OpenStudy (anonymous):

but it is uninteliigible to write -9 -- 4

OpenStudy (anonymous):

yah so therefore because both the numbers are the same sign my lesson said the formula was to " find the sum and keep the sign" i did that and my answer was -13 but wen i chose it and i submitted my test it said i got it wrong, how come?? wat did i do wrong?

OpenStudy (anonymous):

more like there are rules on integers

OpenStudy (anonymous):

yah and i used the ones they gave me in the lesson but it says my answer is wrong

OpenStudy (anonymous):

you're not following it correctly

OpenStudy (anonymous):

tell me how im suppose to follow it

OpenStudy (anonymous):

you owed me 9 bucks, then you 4 bucks of that money you owed

OpenStudy (anonymous):

you returned

OpenStudy (anonymous):

let me type that again you owed me 9 bucks, then you returned 4 bucks of that money that you owed